Suppose you roll a die 3 times. What is the probability that the face 6 will show up at least once?
Suppose we roll a die 2 times. How many outcomes can happen that an even-number face will not show up?
What is the probability that an even number will not show up during the two rolls?
What is the probability that an even number will show up at least once during the two rolls?

Answer by edjones(8007)

A)
(5/6)^3=125/216 probability of no 6 in 3 rolls
1-(5/6)^3=91/216 probability of at least 1 six in 3 rolls.
.
B & C
(1/2)^2=1/4 of the time no even number will show up.
.
D)
1 - 1/4=3/4 of the time an even number will show up.
.
Ed